# MIT-CTP/5902

### v1

@@ D.1 Even $L$
  • Anti-unitary time-reversal $\mathcal{T}$ acts as $\mathcal{T}\chi_{\ell}^{A}\mathcal{T}^{-1}=(-1)^{\ell}\chi_{\ell}^{A}$. The explicit expression for $\mathcal{T}$ depends on the choice of the basis; there is a basis such that $\mathcal{T}$ is just the complex conjugate $\mathcal{K}$ [29].424242 For $N_{f}=1$, the time-reversal operator $\mathcal{T}$ here differs from that in [29] by $(-1)^{F}$.
@@
$\Pi^{2}=1\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (-1)^{F}\Pi=\Pi(-1)^{F}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Pi\,T=(-1)^{\frac{N_{f}(N_{f}-1)}{2}}T^{-1}\,\Pi$ (D.5)
@@
${\tilde{\Pi}}^{2}=(-1)^{\frac{N_{f}(N_{f}-1)}{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ (-1)^{F}{\tilde{\Pi}}=(-1)^{N_{f}}{\tilde{\Pi}}(-1)^{F}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\tilde{\Pi}}T=(-1)^{\frac{N_{f}(N_{f}-1)}{2}}T^{-1}{\tilde{\Pi}}\,$. (D.6)
@@ D.3 A redundant deformation
$\psi_{k,+}=\cos\frac{\theta}{2}d_{-k}+\sin\frac{\theta}{2}d_{-k+N}\,$$\quad$ $\psi_{-k,-}=-\sin\frac{\theta}{2}d_{k}+\cos\frac{\theta}{2}d_{k+N}\,$, (D.25)
@@
$\tan\theta=\lambda^{\prime}\cos(\pi k/N)\,.$
@@
The low-energy Hamiltonian (D.27) shows that this model is gapless. So far, we discussed the even $L$ version of the model. In footnote 45 we mentioned an odd $L$ version of it. Since it differs from the even-$L$ one only by local terms, it remains gapless as well. This is consistent with the constraint from the anomaly of $\tilde{\Pi}$} and $(-1)^{F}$ in Section 2.3.1.
@@
$T(\lambda^{\prime})d_{k}T(\lambda^{\prime})^{-1}=e^{\frac{i\pi k}{N}}\left(\cos\theta d_{k}+\sin\theta d_{k+N}\right)\,.$ (D.29)
@@
$T(\lambda^{\prime})\chi_{\ell}T(\lambda^{\prime})^{-1}$ $=\frac{1}{2N}\sum_{\ell^{\prime}}\left(\sum_{k}e^{i\frac{\pi k(\ell-\ell^{\prime}+1)}{N}}\cos\theta+(-1)^{\ell^{\prime}}\sum_{k}e^{i\frac{\pi k(\ell-\ell^{\prime}+1)}{N}}\sin\theta\right)\chi_{\ell^{\prime}}\,.$ (D.30)
@@ Appendix E Continuum Majorana zero modes on interfaces
${\tilde{\Pi}}\chi_{\text{L}}(t,x){\tilde{\Pi}}^{-1}=\chi_{\text{L}}(-t,-x)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\tilde{\Pi}}\chi_{\text{R}}(t,x){\tilde{\Pi}}^{-1}=\chi_{\text{R}}(-t,-x)\,.$ (E.2)
@@ E.1 Parity-preserving boundary condition
$\chi_{\text{L}}(t,x+2\pi)=\chi_{\text{L}}(t,x)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \chi_{\text{R}}(t,x+2\pi)=\chi_{\text{R}}(t,x)\,.$ (E.3)
@@
[…]
@@
$L_{0}(x)=e^{M(-1+\cos x)}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ R_{0}(x)=-e^{M(-1+\cos x)}\,$, (E.7)
$L_{\pi}(x)=e^{-M(1+\cos x)}\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ R_{\pi}(x)=e^{-M(1+\cos x)}\,.$
@@ E.2 Parity-violating boundary condition
$\chi_{\text{L}}(t,x+2\pi)=-\chi_{\text{L}}(t,x)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \chi_{\text{R}}(t,x+2\pi)=\chi_{\text{R}}(t,x)\,.$ (E.8)
@@
$L(x)=\sinh(M(\pi-|x|))\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ R(x)=-\cosh(M(\pi-|x|))\,$, (E.9)
@@
$L(x)=\sinh(M(1+\cos x))\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ R(x)=-\cosh(M(1+\cos x))\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\pi<x\leq\pi\,.$ (E.10)
@@ F.1 $O(3)$ sigma model
${\cal L}={f\over 2}\partial_{\mu}\vec{n}\cdot\partial^{\mu}\vec{n}+\theta\Gamma\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \vec{n}\cdot\vec{n}=1$ (F.1)
@@
$\mathsf{C}\,\vec{n}(t,x)\,\mathsf{C}^{-1}=-\vec{n}(t,x)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathsf{R}\,\vec{n}(t,x)\,\mathsf{R}^{-1}=\vec{n}(t,-x)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathsf{T}\,\vec{n}(t,x)\,\mathsf{T}^{-1}=-\vec{n}(-t,x)\,.$ (F.2)
@@
$\mathsf{C}\,\Gamma(t,x)\,\mathsf{C}^{-1}=-\Gamma(t,x)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathsf{R}\,\Gamma(t,x)\,\mathsf{R}^{-1}=-\Gamma(t,-x)\,,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathsf{T}\,\Gamma(t,x)\,\mathsf{T}^{-1}=\Gamma(-t,x)\,.$ (F.3)
@@ F.3 Comparing with the lattice
The Hamiltonian (F.10) is in an anti-ferromagnetic phase. In that phase, the lattice operators $\vec{S}_{j}$ do not have a smooth continuum limit. Instead, $(-1)^{j}\vec{S}_{j}$ is smooth. It is natural to identify the continuum limit as [77]
@@
As we said above, the operator on the right-hand side is not the same as the classical field $\vec{n}$. It is a quantum operator with the same quantum numbers. Also, this relation is up to an infinite renormalization factor. We also have [77]

### v2

@@ D.1 Even $L$
  • Anti-unitary time-reversal $\mathcal{T}$ acts as $\mathcal{T}\chi_{\ell}^{A}\mathcal{T}^{-1}=(-1)^{\ell}\chi_{\ell}^{A}$. The explicit expression for $\mathcal{T}$ depends on the choice of the basis; there is a basis such that $\mathcal{T}$ is just the complex conjugate $\mathcal{K}$ [29]. [^43] For $N_{f}=1$, the time-reversal operator $\mathcal{T}$ here differs from that in [29] by $(-1)^{F}$.
@@
$\Pi^{2}=1\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ (-1)^{F}\Pi=\Pi(-1)^{F}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \Pi\,T=(-1)^{\frac{N_{f}(N_{f}-1)}{2}}T^{-1}\,\Pi$ (D.5)
@@
${\tilde{\Pi}}^{2}=(-1)^{\frac{N_{f}(N_{f}-1)}{2}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ (-1)^{F}{\tilde{\Pi}}=(-1)^{N_{f}}{\tilde{\Pi}}(-1)^{F}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ {\tilde{\Pi}}T=(-1)^{\frac{N_{f}(N_{f}-1)}{2}}T^{-1}{\tilde{\Pi}}\,$. (D.6)
@@ D.3 A redundant deformation
$\psi_{k,+}=\cos\frac{\theta_{k}}{2}d_{-k}+\sin\frac{\theta_{k}}{2}d_{-k+N}\,$$\quad$ $\psi_{-k,-}=-\sin\frac{\theta_{k}}{2}d_{k}+\cos\frac{\theta_{k}}{2}d_{k+N}\,$, (D.25)
@@
$\tan\theta_{k}=\lambda^{\prime}\cos(\pi k/N)\,.$
@@
The low-energy Hamiltonian (D.27) shows that this model is gapless. So far, we discussed the even $L$ version of the model. In footnote 46 we mentioned an odd $L$ version of it. Since it differs from the even-$L$ one only by local terms, it remains gapless as well. This is consistent with the constraint from the anomaly of $\tilde{\Pi}$} and $(-1)^{F}$ in Section 2.3.1.
@@
$T(\lambda^{\prime})d_{k}T(\lambda^{\prime})^{-1}=e^{\frac{i\pi k}{N}}\left(\cos\theta_{k}d_{k}+\sin\theta_{k}d_{k+N}\right)\,.$ (D.29)
@@
$T(\lambda^{\prime})\chi_{\ell}T(\lambda^{\prime})^{-1}$ $=\frac{1}{2N}\sum_{\ell^{\prime}}\left(\sum_{k}e^{i\frac{\pi k(\ell-\ell^{\prime}+1)}{N}}\cos\theta_{k}+(-1)^{\ell^{\prime}}\sum_{k}e^{i\frac{\pi k(\ell-\ell^{\prime}+1)}{N}}\sin\theta_{k}\right)\chi_{\ell^{\prime}}\,.$ (D.30)
@@ Appendix E Continuum Majorana zero modes on interfaces
${\tilde{\Pi}}\chi_{\text{L}}(t,x){\tilde{\Pi}}^{-1}=\chi_{\text{L}}(-t,-x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ {\tilde{\Pi}}\chi_{\text{R}}(t,x){\tilde{\Pi}}^{-1}=\chi_{\text{R}}(-t,-x)\,.$ (E.2)
@@ E.1 Parity-preserving boundary condition
$\chi_{\text{L}}(t,x+2\pi)=\chi_{\text{L}}(t,x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \chi_{\text{R}}(t,x+2\pi)=\chi_{\text{R}}(t,x)\,.$ (E.3)
@@
$L^{\prime}(x)=-MR(x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ R^{\prime}(x)=-ML(x)\,$$penalty 10000\ \penalty 10000\ \penalty 10000\$$-\pi<x<0\,$, (E.5)
$L^{\prime}(x)=+MR(x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ R^{\prime}(x)=+ML(x)\,$$penalty 10000\ \penalty 10000\ \penalty 10000\$$0<x\leq\pi\,.$
@@
$L_{0}(x)=e^{M(-1+\cos x)}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ R_{0}(x)=-e^{M(-1+\cos x)}\,$, (E.7)
$L_{\pi}(x)=e^{-M(1+\cos x)}\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ R_{\pi}(x)=e^{-M(1+\cos x)}\,.$
@@ E.2 Parity-violating boundary condition
$\chi_{\text{L}}(t,x+2\pi)=-\chi_{\text{L}}(t,x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \chi_{\text{R}}(t,x+2\pi)=\chi_{\text{R}}(t,x)\,.$ (E.8)
@@
$L(x)=\sinh(M(\pi-|x|))\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ R(x)=-\cosh(M(\pi-|x|))\,$, (E.9)
@@
$L(x)=\sinh(M(1+\cos x))\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ R(x)=-\cosh(M(1+\cos x))\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ -\pi<x\leq\pi\,.$ (E.10)
@@ F.1 $O(3)$ sigma model
${\cal L}={f\over 2}\partial_{\mu}\vec{n}\cdot\partial^{\mu}\vec{n}+\theta\Gamma\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \vec{n}\cdot\vec{n}=1$ (F.1)
@@
$\mathsf{C}\,\vec{n}(t,x)\,\mathsf{C}^{-1}=-\vec{n}(t,x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathsf{R}\,\vec{n}(t,x)\,\mathsf{R}^{-1}=\vec{n}(t,-x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathsf{T}\,\vec{n}(t,x)\,\mathsf{T}^{-1}=-\vec{n}(-t,x)\,.$ (F.2)
@@
$\mathsf{C}\,\Gamma(t,x)\,\mathsf{C}^{-1}=-\Gamma(t,x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathsf{R}\,\Gamma(t,x)\,\mathsf{R}^{-1}=-\Gamma(t,-x)\,,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mathsf{T}\,\Gamma(t,x)\,\mathsf{T}^{-1}=\Gamma(-t,x)\,.$ (F.3)
@@ F.3 Comparing with the lattice
The Hamiltonian (F.10) is in an anti-ferromagnetic phase. In that phase, the lattice operators $\vec{S}_{j}$ do not have a smooth continuum limit. Instead, $(-1)^{j}\vec{S}_{j}$ is smooth. It is natural to identify the continuum limit as [14, 80]
@@
As we said above, the operator on the right-hand side is not the same as the classical field $\vec{n}$. It is a quantum operator with the same quantum numbers. Also, this relation is up to an infinite renormalization factor. We also have [14, 80]